OpenStudy (anonymous):

Solve the system of equations. x + 3y = -1 2x + 2y = 6

OpenStudy (anonymous):

@Luigi0210 @primeralph @thomaster @jim_thompson5910 @jigglypuff314

OpenStudy (anonymous):

You could mutiply the entire first equation by 2, to start out with.

OpenStudy (anonymous):

so which one of these is the answer (-4, 1) (2, -1) (4, -1) (5, -2)

OpenStudy (anonymous):

Alright. When you multiply the first equation by 2, you get 2x+6y=-2 Then you have your other equation: 2x+2y=6. Now treat these two equations as equal to each other. :)

OpenStudy (anonymous):

so it would be b?

OpenStudy (anonymous):

plzzz help

OpenStudy (anonymous):

@primeralph

OpenStudy (primeralph):

I can't just give you the answer.

OpenStudy (primeralph):

Show some effort in solving it.

OpenStudy (anonymous):

can u run me throughgh is what I'm asking

OpenStudy (anonymous):

You'd get 4y=-8. What's your y value then?

OpenStudy (anonymous):

y=-2

OpenStudy (anonymous):

Which one of those choices has y as -2?

OpenStudy (anonymous):

d!

OpenStudy (anonymous):

can u help with another

OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 *(3, 1) *(-1, 3) *(-1, -3) *(1, 3)

OpenStudy (anonymous):

@primeralph

OpenStudy (anonymous):

Nice job on the first question. For the second one, you can subtract y to get 2x=2. What would be your x value? Then which choice has that c value?

OpenStudy (anonymous):

X value, oops.

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

Right on!